1. Costco wants to maintain a 14% mark-up at retail level. If they paid $550 for a necklace, what is the retail price?

2. You are running a sale on pink flamingos: 20% off the retail price of $19.99. 65 are in stock. What is the total markdown amount

3. If fixed costs are $100, price per unit is $10, and variable costs per unit are $6, then the break-even quantity is
Answer #1

1.

It is not mentioned whether the mark-up is on cost or on the selling price. We assume it is on the selling price i.e. retail price.

So,

Markup (amount) = Retail price * Markup %
or, Retail price - Cost = Retail price * Markup %
or, Retail price * (1 - Markup %) = Cost
or, Retail price = Cost / (1 - Markup %) = 550 / (1 - 0.14) = $639.53

2.

Total markdown amount = 20% * Retail price * Volume = 20%*19.99*65 = $259.87

3.

Break-even qty = Fixed cost / (Selling price - Unit variable cost) = 100 / (10 - 6) = 25 units